1. A ball is thrown horizontally from the roof of a building 9.4m tall and lands 9.7m from the base. What was the balls initial speed?

2. You buy a plastic dart gun, and being a clever physics student you decide to do a quick calculation to find its maximum horizontal range. You shoot the gun straight up, and it takes 3.6s for the dart to land back at the barrel. What is the maximum horizontal range of your gun?

3. A stunt driver wants to make his car jump over 8 cars parked side by side below a horizontal ramp.

a. With what minimum speed must he drive off the horizontal ramp? The vertical height of the ramp is 1.5 m above the cars and the horizontal distance he must clear is 22 m.

ANSWER WAS 40 M/S I need help with part b please!

b. If the ramp is now tilted upward, so that "takeoff angle" is 5.0?above the horizontal, what is the new minimum speed? Express your answer using two significant figures.
Answer #1

1. Use x=x0+v0t+.5at^2

9.7=0+V0t+.5(0)t^2 a=0 since there is no acceleration acting in the x direction

And y=y0+v0t+.5at^2 since its thrown horizonally v0 is 0 for the y direction

so 9.4=0+0+.5(9.81)t^2

solve for t and then plug into the first equation 9.7=0+v0t+0 and solve for v0

2. Solve the same way using the different kinematic equations. (below)

x=x0+v0t+.5at^2

v=v0+at

v^2=v0^2+2a(x2-x1)

These equations can be used in the x or y direction with a simple change of variables.

3. this problem is more complicated but the same concepts. Instead of a velocity or force or height being equal to whats given.. you have to find the angle and use cos(angle) and sin(angle) in the equations to get the answer but use the same method you did in part A to solve.